# Shot-noise in strongly-correlated double quantum spin Hall edges

### v2

@@ C.5.1 Conditions on first derivatives at $x=0,x^{\prime},L$
Next, we turn our attention to boundary conditions involving the first derivatives of the propagators, which we obtain by integrating in an $\varepsilon$-ball around the points $x=0,x^{\prime},L$, and taking the limit as $\varepsilon\to 0$. When the differential operators are expanded out, we obtain that 69 yields the following four equations:
@@ C.5.2 Equations at $x=0,L$
From continuity of 119 at $x=0$, we have
@@ C.5.3 Equations at $x=x^{\prime}$
$\bm{\mathbf{G}}^{\text{ret}}(x,x^{\prime},\omega\to 0)\simeq\begin{pmatrix}\frac{K_{\infty}}{2\omega}&0\\ 0&\frac{K_{\infty}}{2\omega}\end{pmatrix}.$ (148)
@@ D.1 Expression for the exact backscattering current
$D^{\alpha\beta}G^{\beta\gamma}=\delta^{\alpha\gamma}\delta^{(2)}(\bm{\mathbf{x}}-\bm{\mathbf{x^{\prime}}})$ (154)
@@ D.2 Expression for the shot noise
The shot noise involves a current-current correlator of the form $\left\langle\partial_{t}\eta(t,x)\partial_{t}\eta(t^{\prime},x^{\prime})\right\rangle$. We may thus consider 150 when $\mathcal{F}=\varphi_{\alpha}$:
$\begin{split}\left\langle\mathcal{F}D^{\gamma\delta}\varphi_{\delta}\right\rangle&=\left\langle i\frac{\delta}{\delta\varphi_{\gamma}}\mathcal{F}-\mathcal{F}\frac{\delta S_{\text{b}}}{\delta\varphi_{\gamma}}\right\rangle\\ &\Rightarrow\left\langle\varphi_{\alpha}D^{\gamma\delta}\varphi_{\delta}\right\rangle=i\delta_{\alpha\gamma}\delta^{(2)}(\bm{\mathbf{x}}^{\prime}-\bm{\mathbf{x}}^{\prime\prime})-\left\langle\varphi_{\alpha}\frac{\delta S_{\text{b}}}{\delta\varphi_{\gamma}}\right\rangle.\end{split}$ (162)
@@
Appendix E Shot noise of the weakly-correlated edge.
In this section, we demonstrate that the shot noise is indeed proportional to the average backscattering current for the backscattering term in the weakly-correlated theory. We have that the total average backscattering current is given by
$\left\langle I_{\text{b}}\right\rangle$ =$\overset{\omega\to 0}{=}-g_{\infty}e\int\mathrm{d}te^{i\omega t}\left\langle\frac{\delta S_{\text{b}}}{\delta\eta_{\text{c}}(t)}\right\rangle$ (165)
@@
$\left\langle I_{\text{b}}\right\rangle=ig_{\infty}e\int\mathrm{d}t\ \lambda_{1}^{2}\left\langle\cos(\eta_{\text{c}}(t)+g_{\infty}eVt)\sin(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\varphi_{\text{n}+},\varphi_{\text{s}+},\varphi_{\text{n}-}](t)\\ -\lambda_{2}^{2}\left\langle\sin(\eta_{\text{c}}(t)+g_{\infty}eVt)\cos(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\varphi_{\text{n}+},\varphi_{\text{s}+},\varphi_{\text{n}-}](t)\\ +\lambda_{3}^{2}\left\langle\cos(\eta_{\text{c}}(t)+g_{\infty}eVt)\sin(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\vartheta_{\text{n}+},\varphi_{\text{s}+},\varphi_{\text{n}-}](t)\\ -\lambda_{4}^{2}\left\langle\sin(\eta_{\text{c}}(t)+g_{\infty}eVt)\cos(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\vartheta_{\text{n}+},\varphi_{\text{s}+},\theta_{\text{n}-}](t)\$ (169)
@@ E.1 Shot noise
$\begin{split}\delta I^{2}&=\frac{(g_{\infty}e)^{2}}{4}\sum_{q=\pm 1}\coth\left(\frac{qg_{\infty}eV}{2k_{\mathrm{B}}T}\right)\int\mathrm{d}t\ e^{iqg_{\infty}eVt}\sum_{m=\pm 1}me^{G(mt)}\mathcal{A}(mt)\\ &\overset{T\to 0}{=}\frac{(g_{\infty}e)^{2}}{4}\sum_{q=\pm 1}q\int\mathrm{d}t\ e^{iqg_{\infty}eVt}\sum_{m=\pm}me^{G(mt)}\mathcal{A}(mt)\\ \end{split}$ (176)
@@
By comparing our expressions for noise and current (eqs. 177 and 172), we see that
@@ F.1 Shot noise
$\begin{split}\delta I^{2}&\overset{\omega\to 0}{=}4(g_{\infty}e)^{2}\int\mathrm{d}t\sum_{m=\pm}\left\langle\sin\left(2\varphi_{\text{c}}(mt)\right)\sin\left(2\varphi_{\text{c}}(0)\right)\right\rangle_{0}\left(\lambda_{2}^{2}+\lambda_{2}^{\prime 2}\mathcal{A}[2\varphi_{\text{s}}](mt)\right)\\ &=(g_{\infty}e)^{2}\sum_{q=\pm}\int\mathrm{d}t\,e^{2iqg_{\infty}eVt}\sum_{m=\pm}e^{2G(mt)}\left(\lambda_{2}^{2}+\lambda_{2}^{\prime 2}\mathcal{A}[2\varphi_{\text{s}}](mt)\right)\\ &=(g_{\infty}e)^{2}\sum_{q=\pm}\coth\left(\frac{2qg_{\infty}eV}{2k_{\mathrm{B}}T}\right)\int\mathrm{d}t\ e^{2iqg_{\infty}eVt}\sum_{m=\pm}me^{2G(mt)}\left(\lambda_{2}^{2}+\lambda_{2}^{\prime 2}\mathcal{A}[2\varphi_{\text{s}}](mt)\right)\\ &\overset{T\to 0}{=}(g_{\infty}e)^{2}\sum_{q=\pm}q\int\mathrm{d}t\ e^{2iqg_{\infty}eVt}\sum_{m=\pm}me^{2G(mt)}\left(\lambda_{2}^{2}+\lambda_{2}^{\prime 2}\mathcal{A}[2\varphi_{\text{s}}](mt)\right)\\ &=2g_{\infty}^{2}e^{2}\int\mathrm{d}t\,e^{2ig_{\infty}et}\sum_{m=\pm}me^{2G(mt)}\left(\lambda_{2}^{2}+\lambda_{2}^{\prime 2}\mathcal{A}[2\varphi_{\text{s}}](mt)\right)\\ \end{split}$ (184)
We can compare this to 183, which establishes the desired result:

### v3

@@ C.5.1 Conditions on first derivatives at $x=0,x^{\prime},L$
Next, we turn our attention to boundary conditions involving the first derivatives of the propagators, which we obtain by integrating in an $\varepsilon$-ball around the points $x=0,x^{\prime},L$, and taking the limit as $\varepsilon\to 0$. When the differential operators are expanded out, we obtain that 72 yields the following four equations:
@@ C.5.2 Equations at $x=0,L$
From continuity of 122 at $x=0$, we have
@@ C.5.3 Equations at $x=x^{\prime}$
$\boldsymbol{\mathbf{G}}^{\text{ret}}(x,x^{\prime},\omega\to 0)\simeq\begin{pmatrix}\frac{K_{\infty}}{2\omega}&0\\ 0&\frac{K_{\infty}}{2\omega}\end{pmatrix}$. (151)
@@ D.1 Expression for the exact backscattering current
$D^{\alpha\beta}G^{\beta\gamma}=\delta^{\alpha\gamma}\delta^{(2)}(\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{x^{\prime}}})$ (157)
@@ D.2 Expression for the shot noise
The shot noise involves a current-current correlator of the form $\left\langle\partial_{t}\eta(t,x)\partial_{t}\eta(t^{\prime},x^{\prime})\right\rangle$. We may thus consider 153 when $\mathcal{F}=\varphi_{\alpha}$:
$\begin{split}\left\langle\mathcal{F}D^{\gamma\delta}\varphi_{\delta}\right\rangle&=\left\langle i\frac{\delta}{\delta\varphi_{\gamma}}\mathcal{F}-\mathcal{F}\frac{\delta S_{\text{b}}}{\delta\varphi_{\gamma}}\right\rangle\\ &\Rightarrow\left\langle\varphi_{\alpha}D^{\gamma\delta}\varphi_{\delta}\right\rangle=i\delta_{\alpha\gamma}\delta^{(2)}(\boldsymbol{\mathbf{x}}^{\prime}-\boldsymbol{\mathbf{x}}^{\prime\prime})-\left\langle\varphi_{\alpha}\frac{\delta S_{\text{b}}}{\delta\varphi_{\gamma}}\right\rangle.\end{split}$ (165)
@@
Appendix E Shot noise of the weakly correlated edge.
In this section, we demonstrate that the shot noise is indeed proportional to the average backscattering current for the backscattering term in the weakly correlated theory. We have that the total average backscattering current is given by
$\left\langle I_{\text{b}}\right\rangle$ $\overset{\omega\to 0}{=}-g_{\infty}e\int\mathrm{d}te^{i\omega t}\left\langle\frac{\delta S_{\text{b}}}{\delta\eta_{\text{c}}(t)}\right\rangle$ (168)
@@
$\left\langle I_{\text{b}}\right\rangle=ig_{\infty}e\int\mathrm{d}t\ \lambda_{1}^{2}\left\langle\cos(\eta_{\text{c}}(t)+g_{\infty}eVt)\sin(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\varphi_{\text{n}+},\varphi_{\text{s}+},\varphi_{\text{n}-}](t)\\ -\lambda_{2}^{2}\left\langle\sin(\eta_{\text{c}}(t)+g_{\infty}eVt)\cos(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\varphi_{\text{n}+},\varphi_{\text{s}+},\varphi_{\text{n}-}](t)\\ +\lambda_{3}^{2}\left\langle\cos(\eta_{\text{c}}(t)+g_{\infty}eVt)\sin(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\vartheta_{\text{n}+},\varphi_{\text{s}+},\varphi_{\text{n}-}](t)\\ -\lambda_{4}^{2}\left\langle\sin(\eta_{\text{c}}(t)+g_{\infty}eVt)\cos(\eta_{\text{c}})\right\rangle_{0}\mathcal{A}[\vartheta_{\text{n}+},\varphi_{\text{s}+},\theta_{\text{n}-}](t)\\$ (172)
@@ E.1 Shot noise
[…]
@@
By comparing our expressions for noise and current (eqs. 180 and 175), we see that
@@ F.1 Shot noise
[…]
We can compare this to 186, which establishes the desired result: